### 05 Mar 2026 v3

# RadarVLM: A Vision-Language Model Approach for Radar Scene Understanding

Abstract
Radar sensors provide reliable perception across adverse weather, lighting, and long-range conditions, yet existing machine learning approaches remain fragmented and task-specific, with each downstream task employing distinct architectures and training objectives. We present RadarVLM, a vision-language framework that learns unified scene-level representations through structured spatial language supervision. Leveraging the CARLA simulator with a realistic radar model, we collect over 800k radar-caption pairs across 110+ hours of simulated driving in diverse scenarios. We make two key contributions: (1) a structured caption framework encoding vehicle distributions in the radar’s native coordinate system, and (2) Spatially-Grounded CLIP (SG-CLIP) objective that replaces binary matching with continuous scene similarity, enabling fine-grained spatial reasoning. We further propose localization-aware evaluation metrics that directly assess spatial accuracy beyond traditional linguistic similarity measures. Validated on generative captioning and vehicle segmentation, SG-CLIP achieves up to 50% relative F1-score improvement over vanilla CLIP and a 21% AP gain on segmentation, demonstrating that language grounding produces spatially structured representations.
Keywords Vision-Language Models $\cdot$ Contrastive Learning $\cdot$ Multimodal Learning $\cdot$ Image-Text Alignment $\cdot$ CLIP
1 Introduction
Figure 1: Overview of the RadarVLM framework. Radar range-angle heatmaps are encoded by a ViT-B/16 vision encoder, while structured spatial captions are processed by a Transformer-based text encoder. Both modalities are projected into a shared embedding space where SG-CLIP aligns representations based on continuous scene similarity. The frozen encoder is subsequently validated via generative captioning (from the CLS token) and vehicle segmentation (from patch tokens).
Autonomous driving systems require robust perception capabilities that operate reliably across diverse environmental conditions. While cameras and LiDAR have driven recent advances, their performance degrades significantly under adverse weather conditions such as rain, fog, and darkness. Radar sensors provide robust, all-weather perception through these adverse weather and lighting conditions [33, 32, 48]. Their ability to directly measure range and velocity makes them indispensable complements to vision-based sensing in autonomous driving. Yet despite these advantages, current machine learning approaches [3, 4, 14, 50, 28] remain fragmented and task-specific. Each downstream task, such as object detection, semantic segmentation, and occupancy prediction, employs distinct input encodings, architectures, and training objectives [16, 7]. This fragmentation results in learned representations that are non-transferable across tasks and fail to generalize to diverse driving scenarios.
The root of this limitation lies in how radar perception has traditionally been framed. Conventional Radar-ML pipelines [8, 18, 34, 45] focus on narrow, isolated objectives with categorical supervision (e.g., bounding boxes for detection or class labels for segmentation). Such supervision lacks the semantic richness needed to capture the complex spatial relationships and contextual cues critical for autonomous driving, such as which lane a vehicle occupies, whether a pedestrian is crossing, or how traffic infrastructure relates to the ego vehicle’s trajectory. What safe driving fundamentally requires is relational spatial reasoning: structured understanding of where objects are, how many there are, and how they are distributed around the ego vehicle. Bounding boxes and class labels cannot encode this.
Language is a natural fit for this gap. A description such as “three vehicles in the right adjacent lane between 10 and 20 m ahead” captures precisely the kind of structured, spatially-grounded information that categorical labels discard. Recent advances in vision-language models [15, 9, 12] have demonstrated that aligning visual representations with natural language produces transferable features that generalize across diverse tasks and unseen categories. Models such as CLIP [38] and BLIP [25] show that language acts as a universal label space, unifying multiple perception objectives under a single representational framework [20, 40]. Crucially, this paradigm is well-suited to radar: unlike images, where appearance varies with lighting and texture, radar heatmaps encode spatial structure directly in range-angle space, which is the same coordinate system in which natural language spatial descriptions are naturally expressed.
Translating this paradigm to radar, however, presents two concrete challenges that prior work has not resolved. First, early radar-language explorations [22, 37] provide only preliminary evidence of alignment, and rely on standard contrastive learning frameworks that treat sample pairs as binary: a matched radar-text pair is positive, all others are negative. This formulation is fundamentally flawed for spatial scene understanding. Two scenes where one has three vehicles in the left lane ahead and the other has two are far more similar to each other than to a scene with no vehicles at all, yet binary labels penalize both equally. This drives the model toward coarse keyword matching rather than fine-grained spatial understanding. Second, pre-training robust radar-language models requires large-scale, well-annotated paired datasets, but real-world radar data collection at scale is expensive and time-consuming, posing a significant barrier.
We introduce RadarVLM (Figure 1), a framework that uses language grounding to teach a radar encoder structured spatial scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [42, 29, 2], enabling large-scale generation of diverse, well-annotated radar-caption pairs across varied autonomous driving scenarios. Our approach makes three key contributions:
  • Structured spatial caption framework: We discretize the radar scene into distance bins and lane-relative angular sectors, which teaches the model not just what objects are present, but where they are which is something that categorical labels do not provide.
  • Spatially-Grounded Contrastive Learning (SG-CLIP): We replace CLIP’s binary matching with a continuous similarity measure based on per-cell vehicle count overlap, enabling fine-grained spatial learning and yielding significant improvements over standard contrastive training.
  • Two-level spatial grounding validation: We validate spatial understanding via generative captioning and patch-level vehicle segmentation, together providing strong evidence that language grounding produces spatially-grounded representations throughout the encoder.
2 Related Works
Vision-Language Models. CLIP [38] demonstrated that contrastive image-text pretraining produces transferable representations, which inspired extensions to dense prediction [27, 53, 9, 15], generative vision-language models [25, 1] which bridge vision encoders with LLMs, and large-scale foundation models [20, 40, 17, 46, 47, 12] which showed that language acts as a universal supervision signal. Why use CLIP? Recent work has shown that the cross-modal alignment behaves like a bag-of-words model [21], failing to preserve attribute-object bindings across modalities. We design our spatial captions to exploit this property by encoding scene semantics as structured distributions.
Radar Perception for Autonomous Driving. Existing Radar ML pipelines fall into three broad categories. Task-specific supervised methods train specialized architectures from scratch for individual perception objectives such as object detection, segmentation, or scene flow [45, 28, 50, 35, 8, 7, 16]. Cross-modal supervised methods exploit external modality supervision to generate pseudo-labels for radar training [3, 34, 8], while SSL approaches use radar-to-radar and radar-to-vision to pretrain radar embeddings [14]. Multi-sensor fusion architectures combine radar with complementary sensors [7, 4]. Across all these paradigms, learned representations remain tied to specific tasks and architectures, and none provide the spatial semantics needed for relational scene understanding.
Contrastive Loss Variants. Several works have extended CLIP’s binary matching with softer supervision, including soft-label relaxations [13], supervised contrastive clustering [19], prototype-based objectives [26], ranking-consistent losses [51], and sigmoid-based or multi-task variants [49, 31]. Closer to our setting, mmCLIP [5] demonstrates the viability of CLIP-style alignment for non-visual sensing by grounding mmWave signals in language. However, none of these address the continuous spatial similarity structure between radar scenes inherent to autonomous driving. Our SG-CLIP objective fills this gap through soft similarity targets derived from per-cell vehicle count overlap.
Radar-Language Alignment. The most closely related works to ours are early explorations of radar-language grounding, this work [37] applies CLIP-style contrastive pretraining between radar spectrograms and text, showing preliminary evidence of improved downstream scene parsing. RadarLLM [22] connects millimeter-wave point cloud sequences to LLMs for human motion understanding. While these works establish the feasibility of radar-language alignment, they rely on generic scene-level descriptions that lack explicit spatial grounding. In contrast, we address both limitations directly, and enables fine-grained spatial learning rather than coarse scene-level matching.
3 Dataset Curation
(a) Lane Distribution
"0-10m": {
"total_vehicles": 3,
"in_lane_front_side": 1,
"right_lane_back_side": 2
"10-20m": {
"total_vehicles": 5,
"in_lane_back_side": 1,
"right_lane_front_side": 3,
"right_lane_back_side": 1
"20-30m": {
"total_vehicles": 4,
"opposing_lane_front": 4
"30-40m": {
"total_vehicles": 2,
"in_lane_front_side": 1,
"right_lane_front_side": 1
"traffic_signs": [],
"walkers": 0
(b) Structured JSON Caption
Figure 2: (a) Lane distribution visualization showing the twelve lane-relative angular sectors used for spatial encoding of vehicles relative to the ego vehicle. (b) Structured JSON scene description from extracted data.
Contrastive VLM pre-training requires large-scale paired data to learn transferable cross-modal representations [38]; however, the scale of real-world annotated radar datasets is not sufficient, motivating our use of simulation. We leverage the CARLA simulator [11], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s radar sensor is highly simplified, we use an open-source implementation [42, 29, 2] which models the propagation behavior of automotive radars very accurately.
Natural language captions: The key challenge in creating effective radar-text pairs lies in designing captions that encode spatially semantics: descriptions that specify where objects are located relative to the ego vehicle rather than merely stating their presence. To address this, we develop a structured encoding framework that partitions the radar scene into distance bins and angular sectors, enabling rich spatial descriptions in radar’s coordinate system.
We discretize the range of 0-40m into four equal bins. Within each distance bin, we categorize vehicles into 12 lane-relative angular sectors based on their positions and heading vectors relative to the ego vehicle (see Figure 2(a)). This ego-centric spatial encoding provides the foundation for generating structured scene descriptions. The vehicle distribution data of every scene is stored in JSON format (see Figure 2(b)).
From the collected JSON representations, we generate natural language captions. Rather than relying on template-based generation, we use LLMs to produce varied descriptions to add diversity among captions. During training, we randomly sample one of the captions for each radar frame, ensuring the model learns from diverse expressions. An example scene from the dataset can be seen in Figure 3(a) with the extracted data in Figure 3(b). The structured JSON caption for the scene is in Figure 2(b) and below is an example of the generated natural language caption:
“ As we are driving, I notice that within the very close range of zero to ten meters from our vehicle, there are three vehicles in total, with one vehicle directly ahead of us in the same lane and two vehicles in the right adjacent lane behind us. Looking a bit further ahead, from ten to twenty meters, I see a total of five vehicles, with three vehicles in the right adjacent lane ahead of us, one vehicle in the right adjacent lane behind us, and one vehicle directly behind us in the same lane. At a moderate distance of twenty to thirty meters, I observe four vehicles in total, all of which are in the opposing lane ahead of us. Further away, from thirty to forty meters, there are two vehicles in total, with one vehicle in the right adjacent lane ahead of us and one vehicle directly ahead of us in the same lane. I also notice that there are no applicable traffic signs around us, and fortunately, there are no walkers on the road. ”
(a) Sensor Views
(b) Extracted Data
Figure 3: Dataset collection overview. (a) Different sensor viewpoints of a traffic example. (b) Extracted data of vehicles in the scene, with lane classification. Dashed circles represent distance radii of 10m, 20m, 30m, and 40m.
Data collection and statistics: We collect diverse driving scenarios across CARLA’s urban, highway, and intersection environments under varying traffic densities. For each radar frame, we extract the complete ground truth state of all actors and filter based on the ego vehicle’s location, retaining only objects within the radar’s sensing range. The traffic scenarios range from sparse (1 to 2 vehicles visible) to dense (10+ vehicles), ensuring the model learns across varying complexity levels. Figure 4(a) shows the distribution of the number of vehicles per scene across the entire dataset, while Figure 4(b) illustrates the lane-wise spatial distribution of collected vehicles. This dataset is, to our knowledge, the first large-scale radar dataset with structured, spatially-grounded natural language descriptions, and we will release it to facilitate future research.
(a) Number of Vehicles per Scene
(b) Distribution of Vehicles across Lanes
Figure 4: Dataset statistics. (a) Distribution of the number of vehicles per scene across the entire dataset. (b) Lane-wise spatial distribution of collected vehicles across the entire dataset. Here, the naming convention follows (Front / Back) - (Lane Position), F stands for front, B stands for back, $L$ stands for Left Lane, IL stands for In Lane, and RL stands for Right Lane.
4 RadarVLM: Architecture and Approach
Our framework, RadarVLM, comprises two key components: a CLIP-based [38, 10] vision encoder for radar heatmaps and a GPT-2-like Transformer [39, 43] text encoder for captions, unified by the Spatially-Grounded Contrastive Learning (SG-CLIP) objective that accounts for continuous scene similarity. While contrastive pretraining aligns radar and text representations at the scene level, it alone cannot confirm that the learned representations are genuinely spatially grounded.
To rigorously validate spatial grounding, we extend RadarVLM on two complementary downstream tasks of generative captioning and segmentation (see Figure 1). Generative captioning interrogates whether the global CLS token encodes sufficient structured semantic information to decode precise vehicle distributions by distance bin and angular sector into accurate natural-language descriptions. Vehicle segmentation directly interrogates the spatial organization of the encoder’s patch-level features, probing whether spatial structure is preserved in the encoder’s internal representations at the pixel level without any linguistic mediation. In both tasks, the vision encoder is kept frozen. A model that succeeds at both has demonstrated spatial grounding at two distinct levels of abstraction, providing strong evidence that SG-CLIP produces representations that are spatially grounded throughout.
4.1 Pre-training Stage
We adopt the pretrained ViT-B/16 encoder from CLIP [38], leveraging its robust visual feature-extraction capabilities trained on large-scale image-text data. The CLS output from the final layer serves as our radar scene representation. For the text encoder, our captions frequently exceed CLIP’s 77-token limit due to their detailed enumeration, so we extended the context window to 400 tokens and trained the encoder from scratch. Both embeddings from radar and text encoders are then projected into a shared 512-dimensional space with MLPs.
4.2 Spatially-Grounded Contrastive Learning
Standard contrastive learning frameworks like CLIP [38] treat sample pairs as binary: a matched image-text pair is positive (label = 1), while all other pairs in the batch are negative (label = 0). This binary formulation is suboptimal because some pairs have partially overlapping spatial configurations of vehicles visible in the radar Range-AoA plot. For instance, two scenes where one contains three vehicles in the left lane ahead and the other contains two are far more similar to each other than to a scene with no vehicles at all. Having binary labels will harshly penalize the model from learning fine-grained spatial distinctions, and instead drive it toward coarse, keyword-based matching.
To address this, we propose the SG-CLIP objective that quantifies scene similarity based on spatial configuration overlap. Let $\mathbf{v}_{i}\in\mathbb{R}_{\geq 0}^{S}$ denote the vector of vehicle counts across all $S$ distance-sector cells for scene $i$. We use these count vectors to compute soft similarity targets between scenes.
Scene dissimilarity: For two scenes $i$ and $j$, we compute the total count discrepancy across all distance bins as:
$d(\mathbf{v}_{i},\mathbf{v}_{j})=\sum_{b=1}^{B_{\text{dist}}}\sum_{s\in\mathcal{S}_{b}}\left|v_{i}^{(b,s)}-v_{j}^{(b,s)}\right|$
where $v_{i}^{(b,s)}$ is the vehicle count in distance bin $b$ and angular sector $s$, and $\mathcal{S}_{b}$ is the set of sectors within bin $b$. Intuitively, $d$ measures the total difference in vehicle counts across all spatial cells between two scenes.
Soft similarity: We convert the dissimilarity into a soft similarity score using a Gaussian kernel:
$s_{ij}=\exp\left(-\alpha\cdot d(\mathbf{v}_{i},\mathbf{v}_{j})^{2}\right)$
where $\alpha$ controls the bandwidth of the kernel. A higher $\alpha$ concentrates similarity mass on nearly identical scenes, approximating the hard binary matching of standard CLIP, whereas a lower $\alpha$ accounts for partially similar scenes, encouraging the model to learn from fine-grained differences. There is an inherent performance tradeoff in varying $\alpha$, described in the next section.
Soft target matrix: We construct a soft target matrix $\mathbf{T}^{\text{soft}}$ for each batch (size $N$) by row-normalizing the pairwise similarities. The normalization ensures a valid probability distribution.
$T_{ij}^{\text{soft}}=\frac{s_{ij}}{\sum_{k=1}^{N}s_{ik}}$
Modified CLIP loss: Given radar embeddings $\mathbf{z}_{r}^{i}$ and text embeddings $\mathbf{z}_{t}^{j}$, we compute pairwise cosine similarities:
$S_{ij}=\frac{\mathbf{z}_{r}^{i}\cdot\mathbf{z}_{t}^{j}}{\|\mathbf{z}_{r}^{i}\|\|\mathbf{z}_{t}^{j}\|}$
The SG-CLIP loss replaces the standard hard cross-entropy with a soft variant:
$\mathcal{L}_{\text{r}\rightarrow\text{t}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}T_{ij}^{\text{soft}}\log\frac{\exp(S_{ij}/\tau)}{\sum_{k=1}^{N}\exp(S_{ik}/\tau)}$
Here $\tau$ is a temperature hyperparameter (we use $\tau=0.07$). The symmetric text-to-radar loss $\mathcal{L}_{\text{t}\rightarrow\text{r}}$ is defined analogously. The final contrastive loss is:
$\mathcal{L}_{\text{SG-CLIP}}=\frac{1}{2}(\mathcal{L}_{\text{r}\rightarrow\text{t}}+\mathcal{L}_{\text{t}\rightarrow\text{r}})$
This formulation provides two key advantages over standard CLIP:
  • Fine-grained spatial learning: The model avoids harsh penalties for placing similar scenes close together in embedding space. Scenes differing by one vehicle receive partial credit rather than being treated as completely different.
  • Scalable supervision: Soft labels are computed automatically from spatial structure without requiring human annotation, allowing to scale naturally to large datasets.
Batch size considerations: The effectiveness of soft contrastive learning depends critically on batch diversity. With soft labels, the model must learn from gradient signals arising from many partially-similar examples with varying degrees of overlap. Empirically, we find that batch sizes upwards of 120 are necessary to provide sufficient diversity for training.
4.3 Generative Captioning
We probe spatial grounding at the level of structured scene semantics by training a lightweight mapping network on top of the frozen encoder’s CLS token. We train a transformer-based mapping network [30] $f_{\theta}$ that projects the CLS embedding into GPT-2’s [39] input space as a set of prefix embeddings. These act as a soft prompt that conditions autoregressive generation, requiring the model to decode the presence of vehicles and their distributions across distance bins and angular sectors.
During training, $f_{\theta}$ is optimized via teacher-forcing [23] to minimize the cross-entropy loss over generated caption tokens:
$\mathcal{L}_{\text{caption}}=-\sum_{t=1}^{T}\log p_{\text{GPT}}(c_{t}\mid\mathbf{p}_{1:k},\,c_{1:t-1})$
where $c_{t}$ is the $t$-th ground-truth caption token, $\mathbf{p}_{1:k}$ are the $k$ prefix embeddings produced by $f_{\theta}$, and $T$ is the caption length. During inference, the radar heatmap is encoded by the frozen vision encoder, passed through $f_{\theta}$ to generate prefix embeddings, without any text input at runtime.
4.4 Vehicle Segmentation
We probe spatial grounding at the patch level by training a lightweight segmentation head on top of the frozen encoder’s patch tokens.
Patch-level feature extraction: During pretraining, the ViT-B/16 encoder produces 196 patch tokens over a $14\times 14$ spatial grid, in addition to the CLS token. While the CLS token aggregates global scene information, the patch tokens retain spatially local features that encode where vehicles appear in the heatmap. We extract these patch tokens from the final transformer layer, apply layer normalization, and reshape them into a feature map $\mathbf{F}\in\mathbb{R}^{768\times 14\times 14}$, which serves as input to the segmentation head (see Figure 1).
Progressive upsampling decoder: To recover pixel-level spatial resolution, we adopt a Progressive UPsampling (PUP) decoder design [52], which has been shown to mitigate the noisy predictions that arise from one-shot large-scale upsampling. The decoder consists of: comprising a convolutional layer with BN and ReLU activation, followed by a $2\times$ bilinear upsampling operation, progressively recovering the full $224\times 224$ resolution. Finally it is followed by a sigmoid activation, which produces a per-pixel probability map $\hat{\mathbf{M}}\in[0,1]^{224\times 224}$.
Ground truth masks: The ground truth segmentation masks are derived from the CARLA simulator’s precise actor positions. For each vehicle in the scene, we project its ground truth location into the radar’s range-angle coordinate frame and construct a Gaussian mask indicating vehicle-occupied pixels. During training, the frozen encoder’s patch features are passed to the segmentation head, keeping the pretrained representations intact while training only the decoder.
Training objective: We train the segmentation head with a combined Soft Dice and Binary Cross-Entropy (BCE) loss, which balances pixel-wise classification accuracy with overlap-based spatial precision. The combined loss is:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{dice}}\cdot\mathcal{L}_{\text{dice}}+\lambda_{\text{bce}}\cdot\mathcal{L}_{\text{bce}}$
The Soft Dice loss directly optimizes the spatial overlap between the predicted probability map and the ground truth mask:
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{p}\hat{M}_{p}\cdot M_{p}+\epsilon}{\sum_{p}\hat{M}_{p}+\sum_{p}M_{p}+\epsilon},\epsilon=10^{-8}$
where $\hat{M}_{p}$ and $M_{p}$ denote the predicted probability and ground truth label at pixel $p$, respectively. The BCE loss provides pixel-wise supervision. We use a weighting of $\lambda_{\text{dice}}=0.6$ and $\lambda_{\text{bce}}=0.4$ in all experiments. The Dice loss is particularly important here, given the class imbalance, wherein the vehicle-occupied pixels are a small fraction of the total scene area.
5 Experiments
Training Setup: All three training stages operate on the same 80/20 train-test split and use the AdamW optimizer, training until validation loss converges. For pretraining, we use a batch size of 160, a cosine learning-rate schedule, and an initial learning rate of $10^{-5}$. The ClipCap captioning decoder is trained with a batch size of 32 and a learning rate of $10^{-4}$, while the segmentation decoder uses a batch size of 512 and a learning rate of $10^{-3}$. All experiments are conducted on a cluster of four NVIDIA A100 GPUs. Wall-clock training time is approximately one day for all tasks.
SG-CLIP produces radar representations that improve both structured scene description and pixel-level spatial localization over standard CLIP-style pre-training. We first qualitatively verify that SG-CLIP concentrates encoder attention on vehicle-occupied regions (Section 5.1), and then quantitatively demonstrate this through: (1) generative captioning assessed with localization-aware metrics, where SG-CLIP achieves up to 50% relative F1-score improvement at long range over vanilla CLIP [38] (Section 5.3), and (2) vehicle segmentation, where SG-CLIP features yield a 5% IoU gain and 21% AP gain over vanilla CLIP [38] and UNet [41, 52] based segmentations (Section 5.4).
5.1 Pre-training Quality: Attention Analysis
Attention rollout [6] recursively multiplies per-layer attention matrices (augmented with the identity for residual connections) across transformer layers, extracting the CLS row of the resulting token-to-token flow matrix as an approximation of cumulative attention to input patches. This provides a direct visualization of which spatial regions the encoder prioritizes.
Figure 5 shows that attention from the CLS token concentrates precisely on spatial regions containing vehicles, with minimal weight allocated to empty sectors. This provides a direct qualitative evidence that SG-CLIP training teaches the encoder to prioritize semantically relevant regions.
Figure 5: Attention rollout from the SG-CLIP pretrained ViT-B/16 encoder, aggregated over the final three transformer layers and overlaid on the radar heatmap. Attention concentrates precisely on vehicle-occupied regions.
5.2 Metrics for Caption Generation
While standard captioning metrics [36, 24, 44] effectively measure n-gram overlap, they are fundamentally ill-suited for assessing spatial accuracy. These metrics treat captions as bags of words or phrases, rewarding lexical similarity without verifying whether spatial information is correct. To address this gap, we adapt precision and recall metrics to directly quantify spatial reasoning accuracy by measuring how precisely the model predicts vehicle positions.
Adapted Precision and Recall: For each distance-sector cell $(b,s)$, let $y$ denote the ground-truth vehicle count and $\hat{y}$ denote the predicted count extracted from the generated caption. We define the True Positive (TP), False Positive (FP), and False Negative (FN) as:
[…]
Note that true negatives (TN) are not considered here, since captions enumerate only the vehicles that are present rather than explicitly predicting their absence in a given sector. To compute precision and recall, we aggregate TP, FP, and FN counts across all test scenes, following the micro-averaging technique, which weights each vehicle equally regardless of scene complexity. The metrics for each distance-sector cell $(b,s)$ are then defined as:
$\text{Precision}_{b,s}=\frac{\text{TP}_{b,s}}{\text{TP}_{b,s}+\text{FP}_{b,s}},\qquad\text{Recall}_{b,s}=\frac{\text{TP}_{b,s}}{\text{TP}_{b,s}+\text{FN}_{b,s}}$
5.2.1 Advantages Over Standard Metrics
  • Spatial grounding: This metric directly assesses whether the model understands where objects are located.
  • Interpretability: Precision and recall have clear semantic meaning in the context of detection, making results easy to interpret.
  • Fine-grained analysis: Cell-wise evaluation enables detailed diagnosis of failure modes (does the model struggle with left vs. right? Near vs. far?).
5.3 Generative Captioning Performance
In this evaluation, the model generates natural language scene descriptions, which are parsed using an LLM to extract structured JSON representations containing vehicle counts per distance-sector cell. These predictions are compared against ground truth using our localization-aware metrics.
(a) Vanilla CLIP
(b) SG-CLIP ($\alpha=1.0$)
Figure 6: Per-cell F1-score distribution across lane-relative angular sectors and distance bins. SG-CLIP variant achieves more uniform and higher F1 scores compared to vanilla CLIP, with the largest improvements at longer ranges.
Method 0–10m 10–20m 20–30m 30–40m
Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$
Vanilla CLIP [38] 0.961 0.926 0.943 0.900 0.817 0.844 0.742 0.669 0.693 0.673 0.529 0.577
Ours ($\alpha=16.0$) 0.983 0.981 0.982 0.929 0.916 0.922 0.849 0.816 0.829 0.870 0.844 0.856
Ours ($\alpha=4.0$) 0.985 0.981 0.983 0.927 0.918 0.922 0.875 0.834 0.852 0.888 0.848 0.867
Ours ($\alpha=1.0$) 0.986 0.985 0.985 0.938 0.930 0.934 0.863 0.858 0.861 0.887 0.848 0.867
Table 1: Localization-aware captioning evaluation across distance bins. Precision, Recall, and F1 are micro-averaged per distance-sector cell. Results are averaged across angular sectors within each distance bin.
Table 1 reveals several key findings. First, replacing binary contrastive labels with graded spatial supervision consistently improves localization accuracy: all SG-CLIP variants outperform vanilla CLIP across every distance bin, with gains increasing at longer ranges. At 30–40 m, SG-CLIP ($\alpha=1.0$) achieves 0.867 F1 versus vanilla CLIP’s 0.577, a 50% relative improvement which shows that soft spatial targets are most beneficial where perceptual signals are weakest.
Second, softer similarity kernels (lower $\alpha$) yield stronger captioning performance. Note that as $\alpha$ increases, the Gaussian kernel concentrates all mass on exact matches, recovering vanilla CLIP’s binary labels. The $\alpha=1.0$ variant achieves the best F1 at three of four distance bins, confirming that distributing the gradient signal across partially similar scenes improves fine-grained spatial reasoning over near-binary matching. Third, while performance degrades with distance for all methods, as expected given weaker vehicle reflections at range, the widening performance gap between SG-CLIP and vanilla CLIP indicates that graded supervision specifically addresses this long-range challenge. Figure 6 visualizes the per-cell F1 distribution, showing that SG-CLIP variants achieve more spatially uniform accuracy across lane sectors compared to vanilla CLIP.
5.4 Vehicle Segmentation
This setup directly tests whether contrastive pre-training, which operates only on the global CLS token, induces spatial structure in local patch representations. Table 2 presents vehicle segmentation results using frozen encoder features. SG-CLIP ($\alpha=4.0$) achieves the strongest overall performance with 0.637 IoU@0.5 and 0.634 AP, outperforming vanilla CLIP by 5% in IoU and 21% in AP.
Method Threshold = 0.5 Peak IoU $\uparrow$ AP $\uparrow$
Prec $\uparrow$ Rec $\uparrow$ IoU $\uparrow$ Dice $\uparrow$
U-Net [41] 0.519 0.927 0.489 0.657 0.489 0.442
Vanilla CLIP [38] 0.820 0.699 0.606 0.755 0.615 0.522
SG-CLIP ($\alpha=1.0$) 0.826 0.717 0.623 0.768 0.631 0.628
SG-CLIP ($\alpha=16.0$) 0.833 0.724 0.625 0.769 0.634 0.631
SG-CLIP ($\alpha=4.0$) 0.848 0.719 0.637 0.778 0.649 0.634
Table 2: Vehicle segmentation results on radar range-angle heatmaps. Precision, Recall and IoU are taken at a constant threshold of 0.5, whereas Peak IoU is the maximum IoU across all confidence thresholds, and AP is the area under Precision-Recall curve.
Several trends stand out. First, all language-pretrained encoders substantially outperform the pre-trained U-Net [41], which is trained on the dataset (0.489 IoU), despite using only a lightweight decoder on frozen features. This confirms that contrastive pre-training with VLMs transfers meaningful spatial structure to patch-level representations, even though it operates only on the global CLS token. Second, the optimal bandwidth ($\alpha=4.0$) differs from the captioning-optimal $\alpha=1.0$, consistent with the different abstraction levels each task probes: segmentation benefits from moderately hard contrastive objectives that sharpen spatial boundaries, while captioning benefits from softer kernels that encourage fine-grained distributional reasoning. Third, the 21% AP improvement from vanilla CLIP to SG-CLIP indicates that graded spatial supervision substantially improves the encoder’s ability to discriminate vehicle-occupied regions across confidence thresholds. Figure 9 confirms this trend, showing SG-CLIP variants dominating across all thresholds.
Figure 7: IoU vs. confidence threshold
Figure 8: Dice vs. confidence threshold
Figure 9: Segmentation performance across confidence thresholds. SG-CLIP variants consistently outperform baselines, with the largest gains at lower thresholds.
6 Conclusion
In this work, we presented RadarVLM, a novel VLM paradigm that shifts radar perception from fragmented, task-specific supervised learning to a unified semantic representation. Our proposed SG-CLIP objective addresses a fundamental limitation in standard contrastive learning by introducing a continuous measure of spatial similarity, allowing to learn nuanced relational dynamics between actors. We demonstrate its efficacy through a large-scale dataset of over 800,000 radar-caption pairs, proving that a single language-grounded encoder can support both generative scene description and discriminative spatial tasks. Furthermore, we argue that this language-mediated supervision provides a robust semantic bridge for sim-to-real transfer, as linguistic spatial relationships remain invariant. Future work will focus on integrating RadarVLM into E2E autonomous driving and validating the generalizability on real-world radar datasets.