### 12 Mar 2026 v4

# RLM: A Vision-Language Model Approach for Radar Scene Understanding

Abstract
Radar sensors enable reliable perception across adverse weather, low-visibility, and long-range conditions, making them critical for robust autonomous driving systems. However, current machine learning approaches for radar remain fragmented, with downstream tasks relying on separate, task‑specific architectures and training objectives. We present Radar Language Model (RLM), a vision-language framework that learns unified scene-level representations through structured spatial language supervision. Leveraging the CARLA simulator with a realistic radar model, we collect over 800k radar-caption pairs across 110+ hours of simulated driving in diverse scenarios. We make two key contributions: (1) a structured caption framework encoding vehicle distributions in the radar’s native coordinate system, and (2) Spatially-Grounded CLIP (SG-CLIP) objective that replaces binary matching with continuous scene similarity, enabling fine-grained spatial reasoning. We further propose localization-aware evaluation metrics. Validated on generative captioning and vehicle segmentation, SG-CLIP consistently outperforms Vanilla CLIP, achieving upto 50% relative F1-score improvement and a 21% AP gain on segmentation, demonstrating that language grounding produces spatially structured representations.
Keywords Vision-Language Models $\cdot$ Contrastive Learning $\cdot$ Multimodal Learning $\cdot$ Image-Text Alignment $\cdot$ CLIP
1 Introduction
Figure 1: Overview of the Radar Language Model (RLM) framework. Radar range-angle heatmaps are encoded by a ViT-B/16 vision encoder, while structured spatial captions are processed by a Transformer-based text encoder. Both modalities are projected into a shared embedding space where SG-CLIP aligns representations based on continuous scene similarity. The frozen encoder is subsequently validated via generative captioning (from the CLS token) and vehicle segmentation (from patch tokens).
Autonomous driving systems require robust perception capabilities that operate reliably across diverse environmental conditions. While cameras and LiDAR have driven recent advances, their performance degrades significantly under adverse weather conditions such as rain, fog, and darkness. Radar sensors provide robust, all-weather perception through these adverse weather and lighting conditions [34, 33, 50]. Their ability to directly measure range and velocity makes them indispensable complements to vision-based sensing in autonomous driving. Yet despite these advantages, current machine learning approaches [3, 4, 15, 52, 29] remain fragmented and task-specific. Each downstream task, such as object detection, semantic segmentation, and occupancy prediction, employs distinct input encodings, architectures, and training objectives [17, 8]. This fragmentation results in learned representations that are non-transferable across tasks and fail to generalize to diverse driving scenarios.
The root of this limitation lies in how radar perception has traditionally been framed. Conventional Radar-ML pipelines [9, 19, 35, 47] focus on narrow, isolated objectives with categorical supervision (e.g., bounding boxes for detection or class labels for segmentation). Such supervision lacks the semantic richness needed to capture the complex spatial relationships and contextual cues critical for autonomous driving, such as which lane a vehicle occupies, whether a pedestrian is crossing, or how traffic infrastructure relates to the ego vehicle’s trajectory. What safe driving fundamentally requires is relational spatial reasoning: structured understanding of where objects are, how many there are, and how they are distributed around the ego vehicle. Bounding boxes and class labels cannot encode this.
Language is a natural fit for this gap. A description such as “three vehicles in the right adjacent lane between 10 and 20 m ahead” captures precisely the kind of structured, spatially-grounded information that categorical labels discard. Recent advances in vision-language models [16, 10, 13] have demonstrated that aligning visual representations with natural language produces transferable features that generalize across diverse tasks and unseen categories. Models such as CLIP [39] and BLIP [26] show that language acts as a universal label space, unifying multiple perception objectives under a single representational framework [21, 41]. Crucially, this paradigm is well-suited to radar: unlike images, where appearance varies with lighting and texture, radar heatmaps encode spatial structure directly in range-angle space, which is the same coordinate system in which natural language spatial descriptions are naturally expressed.
Translating this paradigm to radar, however, presents two concrete challenges that prior work has not resolved. First, early radar-language explorations [23, 38] provide only preliminary evidence of alignment, and rely on standard contrastive learning frameworks that treat sample pairs as binary: a matched radar-text pair is positive, all others are negative. This formulation is fundamentally flawed for spatial scene understanding. Two scenes where one has three vehicles in the left lane ahead and the other has two are far more similar to each other than to a scene with no vehicles at all, yet binary labels penalize both equally. This drives the model toward coarse keyword matching rather than fine-grained spatial understanding. Second, pre-training robust radar-language models requires large-scale, well-annotated paired datasets, but real-world radar data collection at scale is expensive and time-consuming, posing a significant barrier.
We introduce Radar Language Model (RLM) (Figure 1), a framework that uses language grounding to teach a radar encoder structured spatial scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [44, 30, 2], enabling large-scale generation of diverse, well-annotated radar-caption pairs across varied autonomous driving scenarios. Our approach makes three key contributions:
  • Structured spatial caption framework: We discretize the radar scene into distance bins and lane-relative angular sectors, which teaches the model not just what objects are present, but where they are which is something that categorical labels do not provide.
  • Spatially-Grounded Contrastive Learning (SG-CLIP): We replace CLIP’s binary matching with a continuous similarity measure based on per-cell vehicle count overlap, enabling fine-grained spatial learning and yielding significant improvements over standard contrastive training.
  • Two-level spatial grounding validation: We validate spatial understanding via generative captioning and patch-level vehicle segmentation, together providing strong evidence that language grounding produces spatially-grounded representations throughout the encoder.
2 Related Works
Vision-Language Models. CLIP [39] demonstrated that contrastive image-text pretraining produces transferable representations, which inspired extensions to dense prediction [28, 55, 10, 16], generative vision-language models [26, 1] which bridge vision encoders with LLMs, and large-scale foundation models [21, 41, 18, 48, 49, 13] which showed that language acts as a universal supervision signal. Why use CLIP? Recent work has shown that the cross-modal alignment behaves like a bag-of-words model [22], failing to preserve attribute-object bindings across modalities. We design our spatial captions to exploit this property by encoding scene semantics as structured distributions.
Radar Perception for Autonomous Driving. Existing Radar ML pipelines fall into three broad categories. Task-specific supervised methods train specialized architectures from scratch for individual perception objectives such as object detection, segmentation, or scene flow [47, 29, 52, 36, 9, 8, 17]. Cross-modal supervised methods exploit external modality supervision to generate pseudo-labels for radar training [3, 35, 9], while SSL approaches use radar-to-radar and radar-to-vision to pretrain radar embeddings [15]. Multi-sensor fusion architectures combine radar with complementary sensors [8, 4]. Across all these paradigms, learned representations remain tied to specific tasks and architectures, and none of these provide the spatial semantics needed for relational scene understanding.
Contrastive Loss Variants. Several works have extended CLIP’s binary matching with softer supervision, including soft-label relaxations [14], supervised contrastive clustering [20], prototype-based objectives [27], ranking-consistent losses [53], and sigmoid-based or multi-task variants [51, 32]. Closer to our setting, mmCLIP [5] demonstrates the viability of CLIP-style alignment for non-visual sensing by grounding mmWave signals in language. Our work shares the broad motivation of replacing binary targets, but differs in how similarity is defined: prior methods derive soft targets from generic embedding distances, label hierarchies, whereas SG-CLIP constructs similarity directly from per-cell vehicle count overlap in an ego-centric spatial grid, unique to radar driving scenes.
Radar-Language Alignment. The most closely related works to ours are early explorations of radar-language grounding, this work [38] applies CLIP-style contrastive pretraining between radar spectrograms and text, showing preliminary evidence of improved downstream scene parsing. RadarLLM [23] connects millimeter-wave point cloud sequences to LLMs for human motion understanding. While these works establish the feasibility of radar-language alignment, they rely on generic scene-level descriptions that lack explicit spatial grounding. In contrast, we address both limitations directly and enable fine-grained spatial learning rather than coarse scene-level matching.
3 Dataset Curation
(a) Lane Distribution
"0-10m": {
"total_vehicles": 3,
"in_lane_front_side": 1,
"right_lane_back_side": 2
"10-20m": {
"total_vehicles": 5,
"in_lane_back_side": 1,
"right_lane_front_side": 3,
"right_lane_back_side": 1
"20-30m": {
"total_vehicles": 4,
"opposing_lane_front": 4
"30-40m": {
"total_vehicles": 2,
"in_lane_front_side": 1,
"right_lane_front_side": 1
"traffic_signs": [],
"walkers": 0
(b) Structured JSON Caption
Figure 2: (a) Lane distribution visualization showing the twelve lane-relative angular sectors used for spatial encoding of vehicles relative to the ego vehicle. (b) Structured JSON scene description from extracted data.
Contrastive VLM pre-training requires large-scale paired data to learn transferable cross-modal representations [39]; however, the scale of real-world annotated radar datasets is not sufficient, motivating our use of simulation. We leverage the CARLA simulator [12], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s radar sensor is highly simplified, we use an open-source implementation [44, 30, 2] which models the propagation behavior of automotive radars very accurately. We present results on this simulated dataset here; to assess generalization beyond simulation, we additionally experiment with the real-world datasets [4, 43] in the supplementary material.
Natural language captions: The key challenge in creating effective radar-text pairs lies in designing captions that encode spatially semantics: descriptions that specify where objects are located relative to the ego vehicle rather than merely stating their presence. To address this, we develop a structured encoding framework that partitions the radar scene into distance bins and angular sectors, enabling rich spatial descriptions in radar’s coordinate system.
We discretize the range of 0-40m into four equal bins. Within each distance bin, we categorize vehicles into 12 lane-relative angular sectors based on their positions and heading vectors relative to the ego vehicle (see Figure 2(a)). This ego-centric spatial encoding provides the foundation for generating structured scene descriptions. The vehicle distribution data of every scene is stored in JSON format (see Figure 2(b)).
From the collected JSON representations, we generate natural language captions. Rather than relying on template-based generation, we use LLMs to produce varied descriptions to add diversity among captions. During training, we randomly sample one of the captions for each radar frame, ensuring the model learns from diverse expressions. An example scene from the dataset can be seen in Figure 3(a) with the extracted data in Figure 3(b). The structured JSON caption for the scene is in Figure 2(b) and below is an example of the generated natural language caption:
“ As we are driving, I notice that within the very close range of zero to ten meters from our vehicle, there are three vehicles in total, with one vehicle directly ahead of us in the same lane and two vehicles in the right adjacent lane behind us. Looking a bit further ahead, from ten to twenty meters, I see a total of five vehicles, with three vehicles in the right adjacent lane ahead of us, one vehicle in the right adjacent lane behind us, and one vehicle directly behind us in the same lane. At a moderate distance of twenty to thirty meters, I observe four vehicles in total, all of which are in the opposing lane ahead of us. Further away, from thirty to forty meters, there are two vehicles in total, with one vehicle in the right adjacent lane ahead of us and one vehicle directly ahead of us in the same lane. I also notice that there are no applicable traffic signs around us, and fortunately, there are no walkers on the road. ”
(a) Sensor Views
(b) Extracted Data
Figure 3: Dataset collection overview. (a) Different sensor viewpoints of a traffic example. (b) Extracted data of vehicles in the scene, with lane classification. Dashed circles represent distance radii of 10m, 20m, 30m, and 40m.
Data collection and statistics: We collected diverse driving scenarios across CARLA’s urban, highway, and intersection environments under varying traffic densities. For each radar frame, we extract the complete ground truth state of all actors and filter based on the ego vehicle’s location, retaining only objects within the radar’s sensing range. The traffic scenarios range from sparse to dense vehicles, ensuring the model learns across varying complexity levels. Figure 4(a) shows the distribution of the number of vehicles per scene across the entire dataset, while Figure 4(b) illustrates the lane-wise spatial distribution of collected vehicles. This dataset is, to our knowledge, the first large-scale radar dataset with structured, spatially-grounded natural language descriptions, and we will release it to facilitate future research.
(a) Number of Vehicles per Scene
(b) Distribution of Vehicles across Lanes
Figure 4: Dataset statistics. (a) Distribution of the number of vehicles per scene across the entire dataset. (b) Lane-wise spatial distribution of collected vehicles across the entire dataset. Here, the naming convention follows (Front / Back) - (Lane Position), F stands for front, B stands for back, $L$ stands for Left Lane, IL stands for In Lane, and RL stands for Right Lane.
4 RLM: Architecture and Approach
Our framework, RLM, comprises two key components: a CLIP-based [39, 11] vision encoder for radar heatmaps and a GPT-2-like Transformer [40, 45] text encoder for captions, unified by the Spatially-Grounded Contrastive Learning (SG-CLIP) objective that accounts for continuous scene similarity. While contrastive pretraining aligns radar and text representations at the scene level, it alone cannot confirm that the learned representations are genuinely spatially grounded.
To rigorously validate spatial grounding, we extend RLM on two complementary downstream tasks of generative captioning and segmentation (see Figure 1). Generative captioning interrogates whether the global CLS token encodes sufficient structured semantic information to decode precise vehicle distributions by distance bin and angular sector into accurate natural-language descriptions. Vehicle segmentation directly interrogates the spatial organization of the encoder’s patch-level features, probing whether spatial structure is preserved in the encoder’s internal representations at the pixel level without any linguistic mediation. In both tasks, the vision encoder is kept frozen. A model that succeeds at both has demonstrated spatial grounding at two distinct levels of abstraction, providing strong evidence that SG-CLIP produces representations that are spatially grounded throughout.
4.1 Pre-training Stage
We adopt the pretrained ViT-B/16 encoder from CLIP [39], leveraging its robust visual feature-extraction capabilities trained on large-scale image-text data. The CLS output from the final layer serves as our radar scene representation. For the text encoder, our captions frequently exceed CLIP’s 77-token limit due to their detailed enumeration, so we extended the context window to 400 tokens and trained the encoder from scratch. Both embeddings from radar and text encoders are then projected into a shared 512-dimensional space with MLPs.
4.2 Spatially-Grounded Contrastive Learning
Standard contrastive learning frameworks like CLIP [39] treat sample pairs as binary: a matched image-text pair is positive (label = 1), while all other pairs in the batch are negative (label = 0). This binary formulation is suboptimal because some pairs have partially overlapping spatial configurations of vehicles visible in the radar Range-AoA plot. For instance, two scenes where one contains three vehicles in the left lane ahead and the other contains two are far more similar to each other than to a scene with no vehicles at all. Having binary labels will harshly penalize the model from learning fine-grained spatial distinctions, and instead drive it toward coarse, keyword-based matching.
To address this, we propose the SG-CLIP objective that quantifies scene similarity based on spatial configuration overlap. Let $\mathbf{v}_{i}\in\mathbb{R}_{\geq 0}^{S}$ denote the vector of vehicle counts across all $S$ distance-sector cells for scene $i$. We use these count vectors to compute soft similarity targets between scenes.
Scene dissimilarity: For two scenes $i$ and $j$, we compute the total count discrepancy across all distance bins as:
$d(\mathbf{v}_{i},\mathbf{v}_{j})=\sum_{b=1}^{B_{\text{dist}}}\sum_{s\in\mathcal{S}_{b}}\left|v_{i}^{(b,s)}-v_{j}^{(b,s)}\right|$
where $v_{i}^{(b,s)}$ is the vehicle count in distance bin $b$ and angular sector $s$, and $\mathcal{S}_{b}$ is the set of sectors within bin $b$. Intuitively, $d$ measures the total difference in vehicle counts across all spatial cells between two scenes.
Soft similarity: We convert the dissimilarity into a soft similarity score using a Gaussian kernel:
$s_{ij}=\exp\left(-\alpha\cdot d(\mathbf{v}_{i},\mathbf{v}_{j})^{2}\right)$
where $\alpha$ controls the bandwidth of the kernel. A higher $\alpha$ concentrates similarity mass on nearly identical scenes, approximating the hard binary matching of standard CLIP, whereas a lower $\alpha$ accounts for partially similar scenes, encouraging the model to learn from fine-grained differences. There is an inherent performance tradeoff in varying $\alpha$, described in the next section.
Soft target matrix: We construct a soft target matrix $\mathbf{T}^{\text{soft}}$ for each batch (size $N$) by row-normalizing the pairwise similarities. The normalization ensures a valid probability distribution and prevents gradient explosion during training.
$T_{ij}^{\text{soft}}=\frac{s_{ij}}{\sum_{k=1}^{N}s_{ik}}$
Modified CLIP loss: Given radar embeddings $\mathbf{z}_{r}^{i}$ and text embeddings $\mathbf{z}_{t}^{j}$, we compute pairwise cosine similarities:
$S_{ij}=\frac{\mathbf{z}_{r}^{i}\cdot\mathbf{z}_{t}^{j}}{\|\mathbf{z}_{r}^{i}\|\|\mathbf{z}_{t}^{j}\|}$
The SG-CLIP loss replaces the standard hard cross-entropy with a soft variant:
$\mathcal{L}_{\text{r}\rightarrow\text{t}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}T_{ij}^{\text{soft}}\log\frac{\exp(S_{ij}/\tau)}{\sum_{k=1}^{N}\exp(S_{ik}/\tau)}$
Here $\tau$ is a temperature hyperparameter (we use $\tau=0.07$). The symmetric text-to-radar loss $\mathcal{L}_{\text{t}\rightarrow\text{r}}$ is defined analogously. The final contrastive loss is:
$\mathcal{L}_{\text{SG-CLIP}}=\frac{1}{2}(\mathcal{L}_{\text{r}\rightarrow\text{t}}+\mathcal{L}_{\text{t}\rightarrow\text{r}})$
This formulation provides two key advantages over standard CLIP:
  • Fine-grained spatial learning: The model avoids harsh penalties for placing similar scenes close together in embedding space. Scenes differing by one vehicle receive partial credit rather than being treated as completely different.
  • Scalable supervision: Soft labels are computed automatically from spatial structure without requiring human annotation, allowing to scale naturally to large datasets.
Batch size considerations: The effectiveness of soft contrastive learning depends critically on batch diversity. With soft labels, the model must learn from gradient signals arising from many partially-similar examples with varying degrees of overlap. Empirically, we find that batch sizes upwards of 120 are necessary to provide sufficient diversity for training.
4.3 Generative Captioning
We probe spatial grounding at the level of structured scene semantics by training a lightweight mapping network on top of the frozen encoder’s CLS token. We train a transformer-based mapping network [31] $f_{\theta}$ that projects the CLS embedding into GPT-2’s [40] input space as a set of prefix embeddings. These act as a soft prompt that conditions autoregressive generation, requiring the model to decode the presence of vehicles and their distributions across distance bins and angular sectors.
During training, $f_{\theta}$ is optimized via teacher-forcing [24] to minimize the cross-entropy loss over generated caption tokens:
$\mathcal{L}_{\text{caption}}=-\sum_{t=1}^{T}\log p_{\text{GPT}}(c_{t}\mid\mathbf{p}_{1:k},\,c_{1:t-1})$
where $c_{t}$ is the $t$-th ground-truth caption token, $\mathbf{p}_{1:k}$ are the $k$ prefix embeddings produced by $f_{\theta}$, and $T$ is the caption length. During inference, the radar heatmap is encoded by the frozen vision encoder, passed through $f_{\theta}$ to generate prefix embeddings, without any text input at runtime.
4.4 Vehicle Segmentation
We probe spatial grounding at the patch level by training a lightweight segmentation head on top of the frozen encoder’s patch tokens.
Patch-level feature extraction: During pretraining, the ViT-B/16 encoder produces 196 patch tokens over a $14\times 14$ spatial grid, in addition to the CLS token. While the CLS token aggregates global scene information, the patch tokens retain spatially local features that encode where vehicles appear in the heatmap. We extract these patch tokens from the final transformer layer, apply layer normalization, and reshape them into a feature map $\mathbf{F}\in\mathbb{R}^{768\times 14\times 14}$, which serves as input to the segmentation head (see Figure 1).
Progressive upsampling decoder: To recover pixel-level spatial resolution, we adopt a Progressive UPsampling (PUP) decoder design [54], which has been shown to mitigate the noisy predictions that arise from one-shot large-scale upsampling. The decoder consists of: comprising a convolutional layer with BN and ReLU activation, followed by a $2\times$ bilinear upsampling operation, progressively recovering the full $224\times 224$ resolution. Finally it is followed by a sigmoid activation, which produces a per-pixel probability map $\hat{\mathbf{M}}\in[0,1]^{224\times 224}$.
Ground truth masks: The ground truth segmentation masks are derived from the CARLA simulator’s precise actor positions. For each vehicle in the scene, we project its ground truth location into the radar’s range-angle coordinate frame and construct a Gaussian mask indicating vehicle-occupied pixels. During training, the frozen encoder’s patch features are passed to the segmentation head, keeping the pretrained representations intact while training only the decoder.
Training objective: We train the segmentation head with a combined Soft Dice and Binary Cross-Entropy (BCE) loss, which balances pixel-wise classification accuracy with overlap-based spatial precision. The combined loss is:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{dice}}\cdot\mathcal{L}_{\text{dice}}+\lambda_{\text{bce}}\cdot\mathcal{L}_{\text{bce}}$
The Soft Dice loss directly optimizes the spatial overlap between the predicted probability map and the ground truth mask:
$\mathcal{L}_{\text{dice}}=1-\frac{2\sum_{p}\hat{M}_{p}\cdot M_{p}+\epsilon}{\sum_{p}\hat{M}_{p}+\sum_{p}M_{p}+\epsilon},\epsilon=10^{-8}$
where $\hat{M}_{p}$ and $M_{p}$ denote the predicted probability and ground truth label at pixel $p$, respectively. The BCE loss provides pixel-wise supervision. We use a weighting of $\lambda_{\text{dice}}=0.6$ and $\lambda_{\text{bce}}=0.4$ in all experiments. The Dice loss is particularly important here, given the class imbalance, wherein the vehicle-occupied pixels are a small fraction of the total scene area.
5 Experiments
Training Setup: All three training stages operate on the same 80/20 train-test split, use the AdamW optimizer, and training proceeds until validation loss converges. For pretraining, we use a batch size of 160, a cosine learning-rate schedule, and an initial learning rate of $10^{-5}$. The ClipCap captioning decoder is trained with a batch size of 32 and a learning rate of $10^{-4}$, while the segmentation decoder uses a batch size of 512 and a learning rate of $10^{-3}$. All experiments are conducted on a cluster of four NVIDIA A100 GPUs. Wall-clock training time is approximately one day for all tasks.
SG-CLIP produces radar representations that improve both structured scene description and pixel-level spatial localization over standard CLIP-style pre-training. We first qualitatively verify that SG-CLIP concentrates encoder attention on vehicle-occupied regions (Section 5.1), and then quantitatively demonstrate this through: (1) generative captioning assessed with localization-aware metrics, where SG-CLIP consistently improves at long range over vanilla CLIP [39] (Section 5.3), and (2) vehicle segmentation, where SG-CLIP features yield a 5% IoU gain and 21% AP gain over vanilla CLIP [39] and UNet [42, 54] based segmentations (Section 5.4).
5.1 Pre-training Quality: Attention Analysis
Attention rollout [6] recursively multiplies per-layer attention matrices (augmented with the identity for residual connections) across transformer layers, extracting the CLS row of the resulting token-to-token flow matrix as an approximation of cumulative attention to input patches. This provides a direct visualization of which spatial regions the encoder prioritizes.
Figure 5 shows that attention from the CLS token concentrates precisely on spatial regions containing vehicles, with minimal weight allocated to empty sectors. This provides a direct qualitative evidence that SG-CLIP training teaches the encoder to prioritize semantically relevant regions.
Figure 5: Attention rollout from the SG-CLIP pretrained ViT-B/16 encoder, aggregated over the final three transformer layers and overlaid on the radar heatmap. Attention concentrates precisely on vehicle-occupied regions.
5.2 Metrics for Caption Generation
While standard captioning metrics [37, 25, 46] effectively measure n-gram overlap, they are fundamentally ill-suited for assessing spatial accuracy. These metrics treat captions as bags of words or phrases, rewarding lexical similarity without verifying whether spatial information is correct. To address this gap, we adapt precision and recall metrics to directly quantify spatial reasoning accuracy by measuring how precisely the model predicts vehicle positions.
Adapted Precision and Recall: For each distance-sector cell $(b,s)$, let $y$ denote the ground-truth vehicle count and $\hat{y}$ denote the predicted count extracted from the generated caption. We define the True Positive (TP), False Positive (FP), and False Negative (FN) as:
[…]
Note that true negatives (TN) are not considered here, since captions enumerate only the vehicles that are present rather than explicitly predicting their absence in a given sector. To compute precision and recall, we aggregate TP, FP, and FN counts across all test scenes, following the micro-averaging technique, which weights each vehicle equally regardless of scene complexity. The metrics for each distance-sector cell $(b,s)$ are then defined as:
$\text{Precision}_{b,s}=\frac{\text{TP}_{b,s}}{\text{TP}_{b,s}+\text{FP}_{b,s}},\qquad\text{Recall}_{b,s}=\frac{\text{TP}_{b,s}}{\text{TP}_{b,s}+\text{FN}_{b,s}}$
5.2.1 Advantages Over Standard Metrics
  • Spatial grounding: This metric directly assesses whether the model understands where objects are located.
  • Interpretability: Precision and recall have clear semantic meaning in the context of detection, making results easy to interpret.
  • Fine-grained analysis: Cell-wise evaluation enables detailed diagnosis of failure modes (does the model struggle with left vs. right? Near vs. far?).
5.3 Generative Captioning Performance
In this evaluation, the model generates natural language scene descriptions, which are parsed using an LLM to extract structured JSON representations containing vehicle counts per distance-sector cell. These predictions are compared against ground truth using our localization-aware metrics.
(a) Vanilla CLIP
(b) SG-CLIP ($\alpha=1.0$)
Figure 6: Per-cell F1-score distribution across lane-relative angular sectors and distance bins. SG-CLIP variant achieves more uniform and higher F1 scores compared to vanilla CLIP, with the largest improvements at longer ranges.
Method 0–10m 10–20m 20–30m 30–40m
Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ Prec $\uparrow$ Rec $\uparrow$ F1 $\uparrow$
Vanilla CLIP [39] 0.961 0.926 0.943 0.900 0.817 0.844 0.742 0.669 0.693 0.673 0.529 0.577
Ours ($\alpha=16.0$) 0.983 0.981 0.982 0.929 0.916 0.922 0.849 0.816 0.829 0.870 0.844 0.856
Ours ($\alpha=4.0$) 0.985 0.981 0.983 0.927 0.918 0.922 0.875 0.834 0.852 0.888 0.848 0.867
Ours ($\alpha=1.0$) 0.986 0.985 0.985 0.938 0.930 0.934 0.863 0.858 0.861 0.887 0.848 0.867
Table 1: Localization-aware captioning evaluation across distance bins. Precision, Recall, and F1 are micro-averaged per distance-sector cell. Results are averaged across angular sectors within each distance bin.
Table 1 reveals several key findings. First, replacing binary contrastive labels with graded spatial supervision consistently improves localization accuracy: all SG-CLIP variants outperform vanilla CLIP across every distance bin, with gains increasing at longer ranges. At 30–40 m, SG-CLIP ($\alpha=1.0$) achieves 0.867 F1 versus vanilla CLIP’s 0.577, a 50% relative improvement which shows that soft spatial targets are most beneficial where perceptual signals are weakest.
Second, softer similarity kernels (lower $\alpha$) yield stronger captioning performance. Note that as $\alpha$ increases, the Gaussian kernel concentrates all mass on exact matches, recovering vanilla CLIP’s binary labels. The $\alpha=1.0$ variant achieves the best F1 at three of four distance bins, confirming that distributing the gradient signal across partially similar scenes improves fine-grained spatial reasoning over near-binary matching.
Third, while performance degrades with distance for all methods, as expected given weaker vehicle reflections at range, the widening performance gap between SG-CLIP and vanilla CLIP indicates that graded supervision specifically addresses this long-range challenge. Figure 6 visualizes the per-cell F1 distribution, showing that SG-CLIP variants achieve more spatially uniform accuracy across lane sectors compared to vanilla CLIP.
5.4 Vehicle Segmentation
This setup directly tests whether contrastive pre-training, which operates only on the global CLS token, induces spatial structure in local patch representations. Table 2 presents vehicle segmentation results using frozen encoder features. SG-CLIP ($\alpha=4.0$) achieves the strongest overall performance with 0.637 IoU@0.5 and 0.634 AP, outperforming vanilla CLIP by 5% in IoU and 21% in AP.
Method Threshold = 0.5 Peak IoU $\uparrow$ AP $\uparrow$
Prec $\uparrow$ Rec $\uparrow$ IoU $\uparrow$ Dice $\uparrow$
U-Net [42] 0.519 0.927 0.489 0.657 0.489 0.442
Vanilla CLIP [39] 0.820 0.699 0.606 0.755 0.615 0.522
SG-CLIP ($\alpha=1.0$) 0.826 0.717 0.623 0.768 0.631 0.628
SG-CLIP ($\alpha=16.0$) 0.833 0.724 0.625 0.769 0.634 0.631
SG-CLIP ($\alpha=4.0$) 0.848 0.719 0.637 0.778 0.649 0.634
Table 2: Vehicle segmentation results on radar range-angle heatmaps. Precision, Recall and IoU are taken at a constant threshold of 0.5, whereas Peak IoU is the maximum IoU across all confidence thresholds, and AP is the area under Precision-Recall curve.
Several trends stand out. First, all language-pretrained encoders substantially outperform the pre-trained U-Net [42], which is trained on the dataset (0.489 IoU), despite using only a lightweight decoder on frozen features. This confirms that contrastive pre-training with VLMs transfers meaningful spatial structure to patch-level representations, even though it operates only on the global CLS token. Second, the optimal bandwidth ($\alpha=4.0$) differs from the captioning-optimal ($\alpha=1.0$), consistent with the different abstraction levels each task probes: segmentation benefits from moderately hard contrastive objectives that sharpen spatial boundaries, while captioning benefits from softer kernels that encourage fine-grained distributional reasoning. Third, the 21% AP improvement from vanilla CLIP to SG-CLIP indicates that graded spatial supervision substantially improves the encoder’s ability to discriminate vehicle-occupied regions across confidence thresholds. Figure 9 confirms this trend, showing SG-CLIP variants dominating across all thresholds.
Figure 7: IoU vs. confidence threshold
Figure 8: Dice vs. confidence threshold
Figure 9: Segmentation performance across confidence thresholds. SG-CLIP variants consistently outperform baselines, with the largest gains at lower thresholds.
6 Conclusion
In this work, we presented RLM, a novel VLM paradigm that shifts radar perception from fragmented, task-specific supervised learning to a unified semantic representation. Our proposed SG-CLIP objective addresses a fundamental limitation of standard contrastive learning by introducing a continuous measure of spatial similarity, enabling the model to learn nuanced relational dynamics between actors. We demonstrate its efficacy through a large-scale dataset, proving that a single language-grounded encoder can support both generative scene description and discriminative spatial tasks. Furthermore, we argue that this language-mediated supervision provides a robust semantic bridge for sim-to-real transfer, as linguistic spatial relationships remain invariant; preliminary experiments on real-world datasets provide initial supporting evidence. Limitations include a focus on vehicles and the open sim-to-real domain gap, which future work will address alongside integrating RLM into end-to-end autonomous driving systems and validating generalizability on real-world radar datasets.
7 Appendix
7.1 CARLA Dataset
Figure 10: A representative data sample from our CARLA dataset. Top: camera reference view. Bottom left to right: LiDAR bird’s-eye view, radar range-angle heatmap, and the ego-centric spatial grid annotation showing per-sector vehicle counts across four distance bins (0–40 m). This scene depicts a Scenario S3 (VRU crossing) event: a cyclist and motorcycle are visible in the in-lane front sector within 10 m, with additional vehicles in the opposing lane at 20–30 m.
Real-world annotated radar datasets remain scarce and small in scale, making large-scale supervised pre-training infeasible directly from real data. To overcome this, we leverage the C-Shenron radar simulator [44, 30], an open-source, high-fidelity implementation that accurately models the propagation behaviour of automotive radars within the CARLA driving simulator [12]. Crucially, CARLA provides perfect ground-truth spatial annotations for every actor in the scene (as illustrated in Figure 10), something that is prohibitively expensive to obtain at scale in the real world. Our dataset therefore, serves as a structured pre-training foundation: the model learns spatially-grounded radar representations in simulation, which are then transferred and validated on real-world radar data. This sim-to-real strategy is well motivated by the fidelity of C-Shenron, whose ability to closely replicate real automotive radar signatures has been established in prior work [44, 30]; we therefore do not analyse simulator realism further here.
To ensure the pre-trained model learns genuine spatial reasoning rather than memorising superficial patterns, we deliberately collect data across the broadest possible range of driving scenarios, spanning every corner case a radar sensor is likely to encounter in deployment. This includes adversarial traffic conflicts, unsignalized junctions, vulnerable road user crossings, adverse weather, and all lighting conditions from noon to night. Data collection spans all daylight conditions, including morning, afternoon, evening, dusk, and night, following the scenario distribution from [7]. The radar sensor is configured to simulate a 77 GHz Texas Instruments cascaded automotive radar system with 86 virtual antennas, providing realistic Range-Azimuth-Doppler measurements. In total, we collected 879,383 radar-caption frame pairs across distinct CARLA maps and several weather presets, described in detail in the following subsections.
7.1.1 Simulation Environment
We collect data across all eight CARLA maps available in CARLA 0.9.10, spanning residential, urban, highway, and rural environments to ensure broad coverage of driving conditions. Table 3 summarises the road network statistics for each map.
Table 3: CARLA map properties used for data collection.
Map Environment Roads Junctions Max Lanes Length
Town01 Small-scale urban residential 122 12 3 4.2 km
Town02 Very small urban residential 84 8 3 2.0 km
Town03 Mixed urban (roundabout, elevated section) 369 35 10 11.3 km
Town04 Mixed (urban core + highway loop) 276 27 7 10.8 km
Town05 Urban grid with elevated section 323 21 7 9.2 km
Town06 Highway / motorway (Michigan left intersections) 200 23 10 8.5 km
Town07 Rural (narrow winding roads) 277 31 9 5.4 km
Town10HD Dense urban, high-detail city 108 9 8 3.4 km
To maximise scenario diversity and cover corner cases that a radar sensor may encounter, data collection is structured around ten adversarial scenario types drawn from the CARLA Leaderboard benchmark [7]. Each scenario injects a distinct traffic conflict at a trigger waypoint, ensuring the model is exposed to varied spatial configurations of actors beyond routine lane-following. Table 4 summarises each adversarial scenario, its NHTSA pre-crash typology, a brief description of the injected event, and the maps on which it is collected.
Table 4: Adversarial scenario types used for data collection. Town abbreviations: T01–T07 = Town01–Town07, T10 = Town10HD.
ID NHTSA Typology Description Maps
S1 Control Loss / Road Surface Hazard Triggered on any straight road section; debris and steering noise injected. T01–T07, T10
S3 VRU Crossing Without Prior Warning Triggered when ego time-to-arrival to the VRU reaches 10 s. T01–T07, T10
S4 Turning at Intersection: Cyclist Conflict Triggered when ego comes within ${\sim}13$ m of the junction exit on a left or right turn. T01–T07, T10
S7 Signalised Intersection: Red Light Violation Triggered at signalised junction; conflicting lane receives simultaneous green. Subtypes: S7left, S7right, S7opposite. T01–T07, T10
S8 Left Turn: Oncoming Encroachment Triggered only when next route maneuver is a left turn (S8left). T01–T07, T10
S9 Right Turn: Cross-Traffic Conflict Triggered only when next route maneuver is a right turn (S9right). T01–T07, T10
S10 Unsignalised Intersection Triggered at any unsignalised junction; no adversary spawned. T03–T07, T10
In addition to the adversarial scenarios above, we include four route-only configurations (Table 5) that collect pure lane-following data under dense ambient traffic with no injected events. These routes traverse multi-lane highway geometry and are named by the directional maneuver sequence encoded in their waypoint yaw progressions.
Table 5: Route-only lane-following configurations. No adversarial events are injected; ego drives through ambient background traffic only.
ID Maneuver Sequence Description Maps
$l$ Left $\rightarrow$ Left Highway curves with a left-then-left waypoint sequence. T04, T05, T06
lr Left $\rightarrow$ Right Highway curves with a left-then-right waypoint sequence. T03–T06, T10
rl Right $\rightarrow$ Left Adjacent-lane mirror of lr; laterally offset by ${\sim}3.5$ m. T03–T06, T10
$r$ Right $\rightarrow$ Right Adjacent-lane mirror of ll; start points offset by ${\sim}6$ m. T04, T05, T06
Weather diversity. To ensure robustness across all-weather conditions, each data collection episode samples its weather configuration independently at runtime. Seven base presets covering the full precipitation spectrum are combined with six sun altitude levels (night at $80^{\circ}$, twilight at $0^{\circ}$, dawn at $5^{\circ}$, sunset at $15^{\circ}$, morning at $35^{\circ}$, noon at $75^{\circ}$) and eight evenly-spaced sun azimuth angles ($0^{\circ},45^{\circ},\ldots,315^{\circ}$), yielding 336 nominal weather configurations. A Gaussian perturbation ($\sigma=10^{\circ}$) is additionally applied to the sun altitude, making the sampling effectively continuous. Table 6 summarises the base presets; altitude and azimuth overrides are applied on top of each preset at episode initialisation.
Table 6: Base weather presets used during data collection. Sun altitude and azimuth are overridden per episode by the runtime sampler (336 nominal combinations, continuously perturbed).
Preset Cloudiness Precipitation Base Altitude ()
ClearNoon 15 0 75
CloudySunset 80 0 15
WetSunset 20 0 15
MidRainSunset 80 30 15
WetCloudySunset 90 0 15
HardRainNoon 90 60 75
SoftRainSunset 90 15 15
7.1.2 Actor State Extraction and Filtering
For each simulation frame, we extract the complete state of all actors present in the CARLA world. Actors include vehicles, pedestrians, and traffic infrastructure elements such as traffic lights and signs. For each actor, we record: unique identifier, 3D location coordinates, transform (pitch, yaw, roll), velocity vector, acceleration vector, traffic light association status, current and ground-truth traffic light states, traffic sign type, and applicable speed limits. This comprehensive state representation enables us to filter and classify actors based on the ego vehicle’s location in real time.
To focus on perception-relevant objects, we apply distance-based filtering relative to the ego vehicle’s position. Actors located beyond 40 meters from the ego vehicle are excluded, as they fall outside the typical sensing and decision-making range for autonomous driving scenarios. For lane-wise classification, we transform actor positions into the ego vehicle’s local coordinate frame where the ego’s forward direction defines the positive x-axis and the left direction defines the positive y-axis. Each vehicle is classified based on its motion direction relative to the ego vehicle using dot products of forward vectors: vehicles are labeled as "same direction" (dot product >0.5>0.5), "opposing direction" (dot product <0.5<-0.5), or "other" (perpendicular or crossing traffic). Within the "same direction" category, lateral offsets determine lane assignment (in-lane, left-lane, or right-lane) using a configurable lane width threshold of 3 meters. Additionally, each vehicle is classified as being ahead, behind, or directly beside the ego vehicle based on the dot product between the ego’s forward vector and the relative position vector to that vehicle. This ego-centric spatial encoding provides the foundation for generating structured, spatially-grounded scene descriptions.
7.1.3 Natural Language Caption Generation
We generate natural language captions from the structured JSON representations using GPT-4.1 mini. Rather than relying on fixed templates, we prompt the language model to produce varied descriptions that maintain semantic consistency while introducing lexical diversity. The prompting structure consists of two components: a system prompt that instructs the model to adopt the perspective of a driver describing the scene to a passenger, and a user prompt that provides the JSON data along with detailed key-explanation mappings.
System Prompt
The system prompt establishes the generation style and constraints:
"You are an expert at describing traffic scenes from the perspective of a driver talking to a passenger. Only output the final caption. Do not include explanations. Also do not use short word forms with apostrophe, use the full expansion of words only."
This prompt ensures naturalistic descriptions while preventing the model from generating meta-commentary or using contractions, which could introduce inconsistencies during training.
Key-Explanation Mappings
The user prompt includes comprehensive mappings that provide semantic context for each spatial encoding element. These mappings are organized into two categories: distance bins and angular sectors.
Distance Bin Mappings:
  • 0-10m: "Objects within 0 to 10 meters from the ego vehicle. These are very close and may require immediate attention."
  • 10-20m: "Objects within 10 to 20 meters from the ego vehicle. These are close and should be monitored closely."
  • 20-30m: "Objects within 20 to 30 meters from the ego vehicle. These are at a moderate distance and should be noted."
  • 30-40m: "Objects within 30 to 40 meters from the ego vehicle. These are at a safe distance but still relevant."
Angular Sector and Feature Mappings:
  • total_vehicles: "Total number of vehicles in the distance bin."
  • left_lane_front_side: "Vehicles in the left adjacent lane ahead of the ego vehicle."
  • left_side: "Vehicles directly to the left side of the ego vehicle."
  • left_lane_back_side: "Vehicles in the left adjacent lane behind the ego vehicle."
  • in_lane_front_side: "Vehicles directly ahead of the ego vehicle in the same lane."
  • in_lane_back_side: "Vehicles directly behind the ego vehicle in the same lane."
  • right_lane_front_side: "Vehicles in the right adjacent lane ahead of the ego vehicle."
  • right_side: "Vehicles directly to the right side of the ego vehicle."
  • right_lane_back_side: "Vehicles in the right adjacent lane behind the ego vehicle."
  • opposing_lane_front: "Vehicles in the opposing lane ahead of the ego vehicle."
  • opposing_lane_back: "Vehicles in the opposing lane behind the ego vehicle."
  • other_lane_front: "Vehicles in perpendicular or intersecting lanes ahead of the ego vehicle."
  • other_lane_back: "Vehicles in perpendicular or intersecting lanes behind the ego vehicle."
  • applicable_traffic_signs: "Traffic signs that are applicable to the ego vehicle."
  • walkers: "Pedestrians in the scene."
Complete Prompt Example
To illustrate the complete prompting structure, consider a scene where the ego vehicle is on a highway with vehicles in multiple lanes. The structured JSON data is first preprocessed: numerical values are converted to word form (e.g., 3 becomes "three"), and keys within each distance bin are randomly shuffled to prevent the model from learning positional biases. The complete user prompt combines the JSON data with the key explanations:
Here is the JSON data:
  "0-10m": {
    "total_vehicles": "three",
    "in_lane_front_side": "one",
    "right_lane_back_side": "two"
  "10-20m": {
    "total_vehicles": "five",
    "right_lane_front_side": "three",
    "in_lane_back_side": "one",
    "right_lane_back_side": "one"
  "20-30m": {
    "total_vehicles": "four",
    "opposing_lane_front": "four"
  "30-40m": {
    "total_vehicles": "two",
    "in_lane_front_side": "one",
    "right_lane_front_side": "one"
  "applicable_traffic_signs": [],
  "walkers": "zero"
Here is the description of every key in the JSON data:
Distance Bin - 0-10m: Objects within 0 to 10 meters from the ego vehicle. These are very close and may require immediate attention.
[… additional distance bin descriptions …]
- total_vehicles: Total number of vehicles in the distance bin.
- in_lane_front_side: Vehicles directly ahead of the ego vehicle in the same lane.
- right_lane_back_side: Vehicles in the right adjacent lane behind the ego vehicle.
[… additional key descriptions …]
Given this data, generate natural language sentences describing the scene. Only provide the description. Do not explain how you arrived at it. Include all the fields provided even if is empty.
Additionally, we randomly vary whether to include the instruction "Do not add any distance bin numbers" to create captions with and without explicit distance references, further increasing linguistic diversity.
Caption Generation Parameters To enhance caption diversity, we generate multiple caption variants for each radar frame using the following sampling parameters: temperature $T=0.5$ to balance creativity with consistency, top-p nucleus sampling with $p=0.9$ to constrain the probability mass while allowing varied word choices, maximum token length of 400 to accommodate detailed scene descriptions, and number of samples $n=5$ to provide multiple linguistic realizations of the same spatial structure. During training, we randomly sample one caption from this set for each radar frame, ensuring the model learns from diverse expressions of identical spatial configurations while maintaining semantic groundedness.
7.2 Architecture and Training Details
7.2.1 Consolidated Hyperparameter Table
Table 7 summarises the key hyperparameters across all three training stages. All stages use the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\varepsilon=10^{-8}$.
Table 7: Key hyperparameters for all three training stages.
Parameter SG-CLIP ClipCap Segmentation
Optimizer AdamW AdamW AdamW
Weight decay 0.05 0.01 0.01
Learning rate $10^{-5}$ $10^{-5}$ $10^{-4}$
LR schedule Cosine Constant Cosine
Warmup epochs 5 5
Warmup init LR $10^{-7}$ $10^{-6}$
Min LR $10^{-6}$ $10^{-5}$
Gradient clipping 1.0 4.0 4.0
Epochs 100 50 50
Batch size (per GPU) 160 12 512
[…]
Effective batch size 640 72 2048
All three stages share a fixed 80/20 train-test split stratified by vehicle count, with scenes filtered to a maximum of 20 vehicles per frame. Training proceeds for the full epoch budget with the last-epoch checkpoint selected for evaluation; no early stopping is applied. All experiments use TF32 precision with PyTorch DDP across GPUs.
7.2.2 Text Encoder Specification
Since our structured spatial captions frequently enumerate vehicle counts across all four distance bins and twelve angular sectors, they substantially exceed the 77-token context limit of the original CLIP text encoder. We therefore extend the context window to 400 tokens by instantiating a new CLIP model with context_length=400. The pretrained ViT-B/16 vision weights are copied directly; the text encoder is trained from scratch, as the pretrained positional embeddings cannot be reused at the new sequence length.
Table 8 summarises the resulting text encoder architecture. The positional embeddings are learned absolute embeddings of shape $400\times 512$, initialised with $\mathcal{N}(0,\,0.01^{2})$. Attention is causal (upper-triangular mask). The output representation is taken at the EOT (end-of-text) token position, passed through ln_final (LayerNorm, width 512), and projected to the 512-dimensional shared embedding space via a single learned linear projection.
Table 8: Text encoder architecture. The encoder is trained from scratch with an extended context window; all other architectural choices mirror the original CLIP ViT-B/16 text tower.
Parameter Value
Architecture GPT-style causal transformer
Context length 400 tokens
Positional encoding Learned absolute ($400\times 512$)
Vocabulary size 49,408
Transformer width 512
Transformer layers 12
Attention heads 8
MLP inner dim 2,048
MLP activation QuickGELU
Attention mask Upper-Triangular (Causal)
Output token EOT position
Projection head Linear $512\to 512$, no bias
7.2.3 Segmentation Decoder Specification
The PUP decoder receives patch token features from the final layer of the frozen ViT-B/16 encoder. The CLS token is discarded; the remaining 196 patch tokens are layer-normalised (ln_post) and reshaped into a spatial feature map of shape $768\times 14\times 14$, reflecting the $14\times 14$ patch grid of ViT-B/16. The decoder progressively recovers the full $224\times 224$ radar heatmap resolution through four stages of $2\times$ bilinear upsampling interleaved with convolutions, as specified in Table 9. The total upsampling factor of $16\times$ exactly inverts the patch size of the ViT-B/16 backbone. The final $1\times 1$ convolution projects to a single channel, followed by a sigmoid activation to produce a per-pixel vehicle probability map $\hat{M}\in[0,1]^{224\times 224}$.
Table 9: PUP segmentation decoder layer specification. Input: $768\times 14\times 14$ patch features from frozen ViT-B/16. Output: $1\times 224\times 224$ vehicle probability map.
# Layer In Channel Out Channel Kernel Norm Activation Spatial Size
1 Conv2d 768 256 $3\times 3$ BN ReLU $14\times 14$
2 Upsample $2\times$ bilinear $28\times 28$
3 Conv2d 256 256 $3\times 3$ BN ReLU $28\times 28$
4 Upsample $2\times$ bilinear $56\times 56$
5 Conv2d 256 128 $3\times 3$ BN ReLU $56\times 56$
6 Upsample $2\times$ bilinear $112\times 112$
7 Conv2d 128 64 $3\times 3$ BN ReLU $112\times 112$
8 Upsample $2\times$ bilinear $224\times 224$
9 Conv2d 64 1 $1\times 1$ Sigmoid $224\times 224$
Ground truth segmentation masks are constructed by projecting each vehicle’s CARLA world position into the radar range-angle frame and placing a Gaussian blob (circle radius 7 px, $\sigma=2.5$) at the projected location, normalised to $[0,1]$. At evaluation time a ground-truth binarisation threshold of 0.3 is applied to account for the soft Gaussian labels, and prediction thresholds are swept from 0 to 1 in steps of 0.01 to compute threshold-independent metrics (AP, Peak IoU).
7.3 Driving Sequence Demonstrations
We provide video demonstrations of RLM’s spatial understanding capabilities evaluated over continuous driving sequences. Using the C-Shenron [44, 30] autopilot, the ego vehicle navigates complete routes across distinct CARLA towns, each representing a different adversarial scenario from our data collection benchmark. Each video frame presents six synchronized views: the camera reference, the LiDAR bird’s-eye view and radar range-angle heatmap, the ground truth vehicle segmentation mask, the SG-CLIP attention overlay ($\alpha=4.0$), and the predicted segmentation maps from both RLM ($\alpha=4.0$) and the Vanilla CLIP baseline. Representative frames from three of these sequences are shown in Figure 11; the remaining scenario videos are provided in the supplementary video files.
Figure 11: Representative frames from continuous driving sequences across three CARLA scenarios (s4, s7, s9 respectively). Each row shows (left to right): camera view, LiDAR BEV, radar heatmap, Vanilla CLIP predicted segmentation, ground truth segmentation, SG-CLIP attention overlay ($\alpha=4.0$), and RLM predicted segmentation ($\alpha=4.0$).
Across all scenarios, two consistent trends emerge. First, the attention rollout from the SG-CLIP pretrained encoder concentrates tightly on vehicle-occupied regions of the radar heatmap, with negligible activation in empty sectors. This demonstrates that SG-CLIP training produces semantically selective attention even without any pixel-level supervision signal. Second, the RLM segmentation maps closely align with the ground truth Gaussian masks in both vehicle count and spatial placement, whereas the Vanilla CLIP baseline produces diffuse predictions with prominent false-positive activations in vehicle-free regions.
Notably, even in highly cluttered scenes with dense, overlapping vehicle configurations, the individual Gaussian vehicle blobs in RLM’s predictions remain clearly separable, whereas Vanilla CLIP merges nearby vehicles into amorphous activation regions. This contrast is particularly striking given that both models are trained on identical data with identical hyperparameters.
The continuous spatial similarity targets appear to specifically address localisation in adversarial scenarios. However, we do observe a failure mode in the videos: in scenes where multiple vehicles occupy the same distance bin at very similar angular positions, the predicted Gaussian blobs occasionally merge into a single elongated activation, failing to resolve individual vehicles. We attribute this to insufficient batch diversity at the effective batch size used during pretraining, which limits the gradient signal from hard near-duplicate spatial configurations. Increasing the batch size is expected to mitigate this, and we identify it as a limitation of the current work.
7.4 Real-World Data Analysis
To assess generalization beyond simulation, we evaluate RLM on the RADIATE dataset [43], which provides real-world automotive radar data collected under diverse weather conditions. We freeze the SG-CLIP vision encoder and fine-tune only the segmentation decoder on an 80/20 split of RADIATE using identical hyperparameters to the CARLA setting, providing a direct test of whether spatially-grounded representations learned in simulation transfer to real sensor data.
Despite the encoder receiving no real-world radar frames during pre-training, SG-CLIP ($\alpha=4.0$) achieves 0.271 Peak IoU and 0.306 AP on RADIATE, outperforming the Vanilla CLIP baseline (0.207 Peak IoU, 0.217 AP). This relative advantage is preserved consistently across all confidence thresholds (Figure 12), and mirrors the trend observed on our CARLA test set. This suggests that the structured spatial supervision of SG-CLIP encodes representations that generalize beyond the simulation domain.
Table 10: Cross-sensor transfer results on RADIATE. The SG-CLIP encoder is frozen; only the segmentation decoder is fine-tuned. The relative advantage of SG-CLIP over Vanilla CLIP is preserved under real-world distribution shift.
Method Threshold = 0.5 Peak IoU $\uparrow$ AP $\uparrow$
Prec $\uparrow$ Rec $\uparrow$ IoU $\uparrow$ Dice $\uparrow$
Vanilla CLIP [39] 0.352 0.323 0.202 0.337 0.207 0.217
SG-CLIP ($\alpha=4.0$) 0.468 0.380 0.265 0.419 0.271 0.306
(a) IoU vs. confidence threshold
(b) Dice vs. confidence threshold
Figure 12: Segmentation performance on RADIATE across confidence thresholds. SG-CLIP ($\alpha=4.0$) consistently outperforms Vanilla CLIP at all operating points, with the performance gap widening at lower thresholds, consistent with SG-CLIP producing sharper, more localised activations even under real-world distribution shift.
Figure 13: Qualitative segmentation results on RADIATE. Each row shows (left to right): the raw radar heatmap, ground truth segmentation mask, Vanilla CLIP prediction, and RLM ($\alpha=4.0$) prediction.
This trend is corroborated qualitatively in Figure 13. In multi-vehicle scenes (sample 2), RLM produces separable blobs matching ground truth positions while Vanilla CLIP merges nearby returns into broader overlapping activations; RLM also suppresses the false-positive activations Vanilla CLIP exhibits in vehicle-free regions (sample 3). However, in sparse single-vehicle scenes (samples 1 and 4), RLM either misses the target or hallucinates activations at incorrect positions, whereas Vanilla CLIP occasionally fires closer to the correct region. Both models fall short of ground truth localization precision on RADIATE, attributable primarily to the sensor domain gap.
The absolute numbers are lower than our CARLA results. RADIATE uses a Navtech CTS350-X mechanical scanning radar with a 360-degree FOV, whereas our encoder is pre-trained exclusively on range-angle heatmaps from four 77 GHz FMCW radars placed in Front, Back, Left and Right, also giving a 360-degree view. But, these two sensor modalities differ substantially in their polar representation, angular resolution, and clutter characteristics, which constitutes a simultaneous domain shift in both data distribution and sensor geometry. Under these conditions, the performance gap is attributed to sensor mismatch rather than to a limitation of the framework itself: the SG-CLIP objective and architecture require no modification for real-world data, and the relative gains it produces over Vanilla CLIP are fully preserved.
The more fundamental constraint is one of data availability. A large-scale real-world dataset pairing phased-array 77 GHz FMCW radar with per-actor spatial annotations at the scale required for VLM pre-training does not currently exist. Such a dataset would allow the SG-CLIP objective to be applied directly on real sensor data, eliminating the sim-to-real gap. We identify its collection as the most important direction for future work.